OpenStudy (anonymous):

Medals and fan for days! Given that cos(2theta)=3/8, with 2theta terminating in Q4. a) Theta terminates in what quadrant? (just divide by 2?) b) Use an appropriate double angle formula to compute an exact value for cos(theta).

OpenStudy (freckles):

\[\cos(2 \theta)=\frac{3}{8}\] \[\cos(2 \theta)=\cos^2(\theta)-\sin^2(\theta)=1-2 \sin^2(\theta)\] cos(2theta) is given replace cos(2 theta) above with 3/8 and solve the equation for sin(theta)

OpenStudy (freckles):

or better yet since you want cos(theta) look at \[\cos(2 \theta)=\cos^2(\theta)-(1-\cos^2(\theta))=2 \cos^2(\theta)-1 \] \[\cos(2 \theta)=2 \cos^2(\theta)-1 \] this instead

OpenStudy (freckles):

this bottom equation replace the cos(2 theta) with 3/8 and solve for cos(theta)

OpenStudy (freckles):

also if you want to find theta you know 2theta is between 270 and 360 so solve this inequality for theta 270<2 theta <360

OpenStudy (freckles):

now when you solve that one equation for cos you will have two values to choose from use the fact where theta is and what sign cos should be there to pick out which sign you need

OpenStudy (anonymous):

awesome, thanks so much!

OpenStudy (freckles):

do you want me to check your answers?

OpenStudy (anonymous):

believe it's in Q2 so would be neg

OpenStudy (anonymous):

yeah give me one second to compute it

OpenStudy (freckles):

omg you are so great that is so right so far

OpenStudy (anonymous):

-(squareroot11)/4?

OpenStudy (freckles):

that is absolutely right

OpenStudy (anonymous):

thanks for the help magical unicorn!

OpenStudy (freckles):

np lol